In triangle ABC, angle B is 3 times that of angle A, and angle C is 6 times larger than angle A minus 4. what are the measures of the angles
Answer by TimothyLamb(4379)

x = measure of A in degrees
y = measure of B in degrees
z = measure of C in degrees
---
y = 3x
z = 6(x - 4)
z = 6x - 24
x + y + z = 180
---
put the system of linear equations into standard form
---
3x - y = 0
6x - z = 24
x + y + z = 180
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 20.4
y = 61.2
z = 98.4
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---